# A Modified 3D Biconical Outflow Model: Spatial Constraints on AGN-driven Outflows

### v1

@@ Abstract
We present a modified outflow model and its application to constrain ionized outflow properties of active galactic nuclei (AGNs). By adding a rotating disk component to the biconical outflow model of Bae16, we find that models with a rotating disk require faster launching velocities ($\lesssim$ 1500 km s-1) than outflow-only models to be consistent with the observed gas kinematics of local type 2 AGNs. We perform Monte Carlo simulations to reproduce the observed distribution of gas kinematics of a large sample ($\sim$ 39,000), constraining the launching velocity and opening angle. While the launching velocity is moderate for the majority of the local AGNs, the notable cases of 2 - 5 % show strong outflows with $V_{max}\sim 1000-1500$ km s-1. By examining the seeing effect based on the mock integral field unit data, we find that the outflow sizes measured based on velocity widths tend to be overestimated when the angular size of the outflow is comparable to or smaller than the seeing. This result highlights the need for more careful treatments of the seeing effect in the outflow size measurement, yet it still supports the lack of global feedback by gas outflows for local AGNs.
@@
software: astropy (Astropy22), matplotlib (matplotlib), numpy (numpy), scipy (scipy)
@@ 1 Introduction
Accreting supermassive black holes (SMBHs), or active galactic nuclei (AGNs), emit a vast amount of energy through various channels, i.e., radiation, gas outflows, and relativistic jets, which are thought to interact with the interstellar medium (ISM) and eventually affect star formation in host galaxies (e.g., Silk98; Fabian12). Recent theoretical studies commonly implement AGN feedback to reproduce primary observables, i.e., galaxy luminosity function (e.g., Hopkins08; Weinberger17; Wellons23). Nonetheless, the overall role and detailed mechanism of AGN feedback remains unclear in observations (see reviews by Harrison17; Harrison24).
@@
AGN-driven gas outflows are widely considered as a primary candidate for a feedback channel. It is now well-known that these outflows are multiphase (e.g., Fiore17; Fluetsch19; Riffel23), ubiquitous (e.g., Mullaney13; Woo16; Rakshit18), and kpc-scales (e.g., Harrison14; Kim23). In particular, the broad wing component of the [O III] $\lambda 5007$ emission line is frequently utilized to investigate kpc-scale warm ionized gas outflows (Boroson05; Greene05). Large spectral surveys such as Sloan Digital Sky Survey (SDSS) enabled the scrutiny of integrated outflow properties within the central region for several tens of thousands of AGNs and their correlations with host galaxy properties (e.g., Woo17; Woo20).
@@
On the other hand, using spatially resolved spectroscopy, many other studies have focused on the detailed outflow structures for a limited number of targets ($\lesssim$ several tens). One of the primary findings in early studies is that the ionized outflows are well-described by a biconical model with two axisymmetric cones (Crenshaw00a; Crenshaw00b; Veilleux01; Das05; Das06). Subsequent studies took into account the dust extinction by the galactic plane (e.g, Crenshaw10a; Crenshaw10b; Fischer10; Fischer11; Fischer13; Bae16) or a rotating disk (e.g., Muller-Sanchez11) to reproduce the [O III] line profile and diverse narrow line region (NLR) kinematics. Marconcini23 recently developed a new modeling tool for analyzing 3D gas kinematics based on the 3D distribution of numerous gas clouds. By using a sample of nearby AGNs, these studies have provided various constraints on the geometry, extent, and energetics of outflows, which are crucial for quantifying outflow energetics and their impact on the host galaxy evolution.
@@
Despite the importance of estimating the outflow parameters, it is challenging to expand these works to a large sample due to observational limitations. For instance, spatially resolved spectroscopic observations are still expensive to be applied to a large sample. Furthermore, while integral field unit (IFU) surveys such as Mapping Nearby Galaxies at Apache Point Observatory (MaNGA) or Sydney- AAO (Australian Astronomical Observatory) Multi-object Integral-field spectroscopy (SAMI) provide spatially resolved kinematics for $\lesssim$ thousands of AGNs (Wylezalek20; Deconto-Machado22; Gatto24; Oh24), they typically have a seeing-limited resolution (FWHM $\sim$ 1″- 2″), which is often insufficient to resolve outflow structures at moderate redshifts. For example, a few kpc scales of outflows at z $\sim$ 0.1 (1 kpc $\sim$ 0.5″) are barely resolved even under the best seeing conditions on the ground. While space telescopes such as Hubble Space Telescope (HST) and James Webb Space Telescope (JWST) or adaptive optics (AO) can improve the spatial resolution down to a diffraction-limit (FWHM $\sim$ 0.1″), the detailed geometry and kinematics can only be resolved in rare cases such as nearby AGNs (e.g., Fischer13; Polack24) or galactic scale outflows (e.g., Vayner21a; Wylezalek22; Cresci23).
@@
Nevertheless, detailed outflow models and simulated mock data can be efficiently utilized to overcome the limited spatial resolution and constrain the physical parameters of the outflows, which are not directly measurable. Bae16 constructed a 3D biconical outflow model combined with a simple extinction by a thin dust plane and investigated the effect of each parameter on the gas kinematics. Their main findings are as follows: (1) velocity dispersion is higher for larger intrinsic outflow velocity or bicone inclination; (2) velocity offset increases with higher dust extinction; (3) a maximum velocity of 500 - 1000 km s-1can explain most of the observed gas kinematics of local SDSS type 2 AGNs from Woo16; and (4) higher intrinsic velocities tend to have wider bicone opening angles. However, this model did not account for rotating disk emission and seeing convolution, which are critical for reproducing 2D gas kinematics.
@@
2 Modified 3D Biconical Outflow Model
@@
We summarize the original outflow model from Bae16 in Section 2.1 and demonstrate three major modifications in Section 2.2, 2.3, and 2.4, respectively.
2.1 3D Biconical Outflow Model
@@
Bae16 presented a 3D outflow model adopted from Crenshaw10a; Crenshaw10b, which consists of two axisymmetric outflow cones and a simple extinction factor by a thin galactic dust plane (see Figure 1). Once a set of 11 parameters is given, the model constructs 3D distributions of flux and velocity. By collapsing the 3D model along the line-of-sight (LOS), 2D distributions of flux, velocity, and velocity dispersion are generated. Bae16 also calculated flux-weighted velocity and velocity dispersion using the 2D maps for a direct comparison with the observed [O III] velocity and velocity dispersion (VVD) distribution of SDSS type 2 AGNs from Woo16.
@@
The input parameters define the bicone structure, 3D distributions of flux and kinematics, and the orientations of the bicone and the dust plane. For the bicone structure, three parameters determine the outflow size ($R_{out}$) and inner and outer half opening angles ($\theta_{in}$ and $\theta_{out}$). As illustrated by the dark gray region in Figure 1, we model gas outflows using a hollow bicone. For simplicity, we set all models in this study to have an outer half opening angle that is 20 larger than the inner one ($\theta_{out}=\theta_{in}+20^{\circ}$).
@@
The flux distribution is determined by e-folding radius ($\tau$) and dust extinction factor ($A$). Since we only focus on relative flux distributions, we adopt an arbitrary central flux value ($f_{0}$). We assume that the intrinsic 3D flux distribution follows an exponential form based on the distance from the center ($R$). Thus, the intrinsic 3D flux distribution can be written as $f(R)=f_{0}e^{-R/\tau}$. Following Bae16, we assume the e-folding radius is 20% of the outflow size ($\tau=R_{out}/5$), which gives approximately two orders of magnitude fainter flux at the outflow edge than the center. The dust extinction factor is multiplied when collapsing the 3D flux distributions to 2D flux maps. If a voxel is located behind the galactic dust plane along the LOS, its flux will be dimmed by the extinction factor in the 2D map. The opening angles and the relative inclination compared to the galactic plane decide which voxels are subject to dust extinction.
@@
The 3D velocity distributions are determined by two parameters, i.e., maximum velocity ($V_{max}$) and velocity profile. There are three options for the velocity profile as a function of the distance from the center ($R$): (1) a linear increase case ($v(R)=V_{max}(R/R_{out})$); (2) a linear decrease case ($v(R)=V_{max}(1-R/R_{out})$); and (3) a constant velocity ($v(R)=V_{max}$). In all three cases, the velocity is directed radially outward from the center. Following Bae16, we adopt the linear decrease case for all models in this study. Note that the maximum velocity can be regarded as a launching velocity, as it represents the intrinsic outflow velocity at the center. This 3D velocity distribution is then collapsed to the 2D velocity and velocity dispersion maps in a flux-weighted manner as follows:
@@
Lastly, the orientations of the model are determined by position angles (PAcone and PAdisk) and inclinations ($i_{cone}$ and $i_{disk}$) of the bicone and the dust plane (or disk in Section 2.2). The orientations of the bicone and the dust plane are independent of each other, allowing them to rotate freely on the sky plane. However, for simplicity, we fix the PAs of the bicone and dust plane to 0 and 90, respectively, meaning that the bicone axis and the major axis of the plane extend along the north-south and east-west directions, respectively. This is because our mock data analysis in Section 3 focuses on 3″-extracted spectra and radial profiles, which are little affected by PAs on the sky plane.
@@
Using this biconical model, Bae16 tested the impact of each parameter on the velocity offset and dispersion, and the line profile of the [O III] emission line. Nevertheless, the model lacked several components crucial for comparison with observational data. For instance, it did not consider rotating gas in a galactic disk, which is thought to mainly contribute to the narrow/core component of the typical [O III] line profile. Furthermore, to accurately compare the spatial properties of the simulated outflows with observations, the seeing convolution should be taken into account. Therefore, we add a rotating gas component (Section 2.2) and the seeing effect (Section 2.3) to the Bae16 model. In addition, to improve mock data compatibility, we switch from a coordinate system normalized by the outflow size to a physical scale in arcsec units based on the redshift and outflow size parameters (Section 2.4).
@@
Figure 1: A schematic diagram of a biconical model example of Bae16. SMBH is located at the center of this plot, and the dust plane is shown as a red line between two cones. The dark gray area represents where the outflowing gas exists (i.e., between inner and outer opening angles), while the light gray area shows a hollow region. The LOS direction is shown as a black arrow on the right.
2.2 Rotating Component
@@
Although the outflow is a primary kinematic component in many AGNs, the gas kinematics in the NLR are more complicated due to the presence of other effects. Among them, gas following the gravitational potential of the host galaxy is often observed as a narrow emission line that has similar kinematics to stellar absorption lines ($\sigma_{[\mathrm{OIII}],narrow}\sim\sigma_{*}$ Nelson96; Greene05). For example, if there is no non-gravitational effect, i.e., outflows, the gas kinematics are dominated by a rotating disk component (e.g., Luo19). Given that a majority of AGNs in the VVD diagram (see Figure 4(a)) are located at relatively small [O III] velocity offset and dispersion, the disk component should be included to properly reproduce the observed distribution of the gas kinematics. Even for AGNs with strong outflows, a rotational pattern is often observed in IFU data (Kang18; Kim23). Thus, rotation due to the host galaxy gravity is required to properly reproduce the demography of the gas kinematics as well as the characteristics in the spatially resolved kinematics.
@@
To generate disk models, we make use of the GalPak3D package (Galaxy Parameters and Kinematics, Bouche15), which was originally developed to model the morphokinematics of 3D observational data. Since our goal is to investigate how the rotating gas component influences the total gas kinematics, we employ only the basic models generated by the package without adjusting details. Consequently, all disk models in this study are based on the default models of GalPak3D. The radial and perpendicular flux profiles are set to exponential ($n=1$) and Gaussian, respectively, and the rotation curve follows a hyperbolic tan (tanh) profile.
@@
We regulate two flux parameters and three kinematic parameters of the disk model. The inclination and PA of the disk are assumed to be the same as those of the dust plane in the outflow model. For the flux, the total disk flux within the SDSS fiber size (3″) is first determined by the disk-to-cone flux ratio parameter ($f_{disk}/f_{cone}$) and the integrated flux of the outflow. We then determine the disk flux distribution by assuming an exponential disk with a given half-light radius parameter ($r_{1/2}$).
@@
where $r$ is the radius on the galactic plane. In addition, we adjust the intrinsic velocity dispersion ($\sigma_{0}$), which corresponds to the dynamical hotness of the disk.
@@
2.3 Seeing Effect
The beam smearing effect by atmospheric seeing has been pointed out as a significant obstacle in spatially resolved studies of AGNs (e.g., Husemann16; Villar-Martin16; Singha22). Given that the seeing effect is more pronounced at higher redshifts, we need to include it in our model to account for outflows in a majority of AGN host galaxies.

### v2

@@ Abstract
We present a modified outflow model and its application to constrain ionized outflow properties of active galactic nuclei (AGNs). By adding a rotating disk component to the biconical outflow model of Bae and Woo (2016), we find that models with a rotating disk require faster launching velocities ($\lesssim$ 1500 km s-1) than outflow-only models to be consistent with the observed gas kinematics of local type 2 AGNs. We perform Monte Carlo simulations to reproduce the observed distribution of gas kinematics of a large sample ($\sim$ 39,000), constraining the launching velocity and opening angle. While the launching velocity is moderate for the majority of the local AGNs, the notable cases of 2 - 5 % show strong outflows with $V_{max}\sim 1000-1500$ km s-1. By examining the seeing effect based on the mock integral field unit data, we find that the outflow sizes measured based on velocity widths tend to be overestimated when the angular size of the outflow is comparable to or smaller than the seeing. This result highlights the need for more careful treatments of the seeing effect in the outflow size measurement, yet it still supports the lack of global feedback by gas outflows for local AGNs.
@@
software: astropy (Astropy Collaboration et al., 2022), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020)
@@ I Introduction
Accreting supermassive black holes, or active galactic nuclei (AGNs), emit a vast amount of energy through various channels, i.e., radiation, gas outflows, and relativistic jets, which are thought to interact with the interstellar medium (ISM) and eventually affect star formation in host galaxies (e.g., Silk and Rees, 1998; Fabian, 2012). Recent theoretical studies commonly implement AGN feedback to reproduce primary observables, i.e., galaxy luminosity function (e.g., Hopkins et al., 2008; Weinberger et al., 2017; Wellons et al., 2023). Nonetheless, the overall role and detailed mechanism of AGN feedback remain unclear in observations (see reviews by Harrison, 2017; Harrison and Ramos Almeida, 2024).
@@
AGN-driven gas outflows are widely considered as one of the primary candidates for a feedback channel. It is now well-known that these outflows are multiphase (e.g., Fiore et al., 2017; Fluetsch et al., 2019; Riffel et al., 2023), ubiquitous (e.g., Mullaney et al., 2013; Woo et al., 2016; Rakshit and Woo, 2018), and kpc-scales (e.g., Harrison et al., 2014; Kim et al., 2023). In particular, the broad wing component of the [O III] $\lambda 5007$ emission line is frequently utilized to investigate kpc-scale warm ionized gas outflows (Boroson, 2005; Greene and Ho, 2005). Large spectral surveys such as Sloan Digital Sky Survey (SDSS) enabled the scrutiny of integrated outflow properties within the central region for several tens of thousands of AGNs and their correlations with host galaxy properties (e.g., Woo et al., 2017, 2020).
@@
On the other hand, using spatially resolved spectroscopy, many other studies have focused on the detailed outflow structures for a limited number of targets ($\lesssim$ several tens). One of the primary findings in early studies is that the ionized outflows are well-described by a biconical model with two axisymmetric cones (Crenshaw and Kraemer, 2000; Crenshaw et al., 2000; Veilleux et al., 2001; Das et al., 2005, 2006). Subsequent studies took into account the dust extinction by the galactic plane (e.g, Crenshaw et al., 2010a, b; Fischer et al., 2010, 2011, 2013; Bae and Woo, 2016) or a rotating disk (e.g., Müller-Sánchez et al., 2011) to reproduce the [O III] line profile and diverse narrow line region (NLR) kinematics. Marconcini et al. (2023) recently developed a new modeling tool for analyzing 3D gas kinematics based on the 3D distribution of numerous gas clouds. By using a sample of nearby AGNs, these studies have provided various constraints on the geometry, extent, and energetics of outflows, which are crucial for quantifying outflow energetics and their impact on the host galaxy evolution.
@@
Despite the importance of estimating the outflow parameters, it is challenging to expand these works to a large sample due to observational limitations. For instance, spatially resolved spectroscopic observations are still too expensive to be applied to a large sample. Furthermore, while integral field unit (IFU) surveys such as MaNGA or SAMI provide spatially resolved kinematics for $\lesssim$ thousands of AGNs (Wylezalek et al., 2020; Deconto-Machado et al., 2022; Gatto et al., 2024; Oh et al., 2024), they typically have a seeing-limited resolution (FWHM $\sim$ 1″- 2″), which is often insufficient to resolve outflow structures at moderate redshifts. For example, a few kpc scales of outflows at z $\sim$ 0.1 (1 kpc $\sim$ 0.5″) are barely resolved even under the best seeing conditions on the ground. While space telescopes such as Hubble Space Telescope (HST) and James Webb Space Telescope (JWST) or adaptive optics (AO) can improve the spatial resolution down to a diffraction-limit (FWHM $\sim$ 0.1″), the detailed geometry and kinematics can only be resolved in rare cases such as nearby AGNs (e.g., Fischer et al., 2013; Polack et al., 2024) or galactic scale outflows (e.g., Vayner et al., 2021; Wylezalek et al., 2022; Cresci et al., 2023).
@@
Nevertheless, detailed outflow models and simulated mock data can be efficiently utilized to overcome the limited spatial resolution and constrain the physical parameters of the outflows, which are not directly measurable. Bae and Woo (2016) constructed a 3D biconical outflow model combined with a simple extinction by a thin dust plane and investigated the effect of each parameter on the gas kinematics. Their main findings are as follows: (1) velocity dispersion is higher for larger intrinsic outflow velocity or bicone inclination; (2) velocity offset with respect to systemic velocity increases with higher dust extinction; (3) a maximum velocity of 500 - 1000 km s-1 can explain most of the observed gas kinematics of local SDSS type 2 AGNs from Woo et al. (2016); and (4) higher intrinsic velocities tend to have wider bicone opening angles. However, this model did not account for rotating disk emission and seeing convolution, which are critical for reproducing 2D gas kinematics.
@@
II Modified 3D Biconical Outflow Model
@@
We summarize the original outflow model from Bae and Woo (2016) in Section II.1 and demonstrate three major modifications in Section II.2, II.3, and II.4, respectively.
II.1 3D Biconical Outflow Model
@@
Bae and Woo (2016) presented a 3D outflow model adopted from Crenshaw et al. (2010a, b), which consists of two axisymmetric outflow cones and a simple extinction factor by a thin galactic dust plane (see Figure 1). Once a set of 11 parameters is given, the model constructs 3D distributions of flux and velocity. By collapsing the 3D model along the line-of-sight (LOS), 2D distributions of flux, velocity, and velocity dispersion are generated. Bae and Woo (2016) also calculated flux-weighted velocity and velocity dispersion using the 2D maps for a direct comparison with the observed [O III] velocity and velocity dispersion (VVD) distribution of SDSS type 2 AGNs from Woo et al. (2016).
@@
The input parameters define the bicone structure, 3D distributions of flux and kinematics, and the orientations of the bicone and the dust plane. For the bicone structure, three parameters determine the outflow size ($R_{out}$) and inner and outer half opening angles ($\theta_{in}$ and $\theta_{out}$). As illustrated by the dark gray region in Figure 1, we model gas outflows using a hollow bicone. For simplicity, we fix the difference between the outer and inner half opening angles to 20 ($\theta_{out}=\theta_{in}+20^{\circ}$). While the opening angle differences estimated by Fischer et al. (2013) are mostly in the 5–20 range, they adopted the minimum opening angle difference required to match the observations, meaning that the larger difference is also plausible. Thus, we adopted 20 as a representative value from observations. We also note that the inner opening angle has an opposite, but less significant effect on the integrated kinematics compared to the outer opening angle (Bae and Woo, 2016).
@@
The flux distribution is determined by $e$-folding radius ($\tau$), i.e., a radius where the flux decreases by a factor of $e$, and dust extinction factor ($A$). Since we only focus on relative flux distributions, we adopt an arbitrary central flux value ($f_{0}$). We assume that the intrinsic 3D flux distribution follows an exponential form based on the distance from the center ($R$). Thus, the intrinsic 3D flux distribution can be written as $f(R)=f_{0}e^{-R/\tau}$. Following Bae and Woo (2016), we assume the $e$-folding radius is 20% of the outflow size ($\tau=R_{out}/5$), which gives approximately two orders of magnitude fainter flux at the outflow edge than at the center. The dust extinction factor is multiplied when collapsing the 3D flux distributions to 2D flux maps. If a voxel is located behind the galactic dust plane along the LOS, its flux will be dimmed by the extinction factor in the 2D map. The opening angles and the relative inclination compared to the galactic plane decide which voxels are subject to dust extinction.
@@
The 3D velocity distributions are determined by two parameters, i.e., maximum velocity ($V_{max}$) and velocity profile. There are three options for the velocity profile as a function of the distance from the center ($R$): (1) a linear increase case ($v(R)=V_{max}(R/R_{out})$); (2) a linear decrease case ($v(R)=V_{max}(1-R/R_{out})$); and (3) a constant velocity ($v(R)=V_{max}$). In all three cases, the velocity is directed radially outward from the center. Following Bae and Woo (2016), we adopt the linear decrease case for all models in this study. Note that the maximum velocity can be regarded as a launching velocity, as it represents the intrinsic outflow velocity at the center. This 3D velocity distribution is then collapsed to the 2D velocity and velocity dispersion maps in a flux-weighted manner as follows:
@@
Lastly, the orientations of the model are determined by position angles (PAs; PAcone and PAdisk) and inclinations ($i_{cone}$ and $i_{disk}$) of the bicone and the dust plane (or disk in Section II.2). The orientations of the bicone and the dust plane are independent of each other, allowing them to rotate freely on the sky plane. However, for simplicity, we fix the PAs of the bicone and dust plane to 0 and 90, respectively, meaning that the bicone axis and the major axis of the plane extend along the north-south and east-west directions, respectively. Since our mock data analysis focuses on 3″-extracted kinematics and radial profiles of velocity widths, the effect of PAs is negligible.
Note that we do not consider various relative PA values between the disk and bicone for simplicity. We expect that the relative PA has no significant effect on mock SDSS data since we integrate all pixels within the 3″-aperture. In addition, velocity widths in mock IFU analysis are dominated by the bicone component in cases of strong outflows, thereby reducing the impact of the rotational orientation of the disk on the kinematics. The effect of relative PAs on mock IFU data will be investigated in future work.
@@
Using this biconical model, Bae and Woo (2016) tested the impact of each parameter on the velocity offset and dispersion, and the line profile of the [O III] emission line. Nevertheless, the model lacked several components crucial for comparison with observational data. For instance, it did not consider rotating gas in a galactic disk, which is thought to mainly contribute to the narrow/core component of the typical [O III] line profile. Furthermore, to accurately compare the spatial properties of the simulated outflows with observations, the seeing convolution should be taken into account. Therefore, we add a rotating gas component (Section II.2) and the seeing effect (Section II.3) to the Bae and Woo (2016) model. In addition, to improve mock data compatibility, we switch from a coordinate system normalized by the outflow size to a physical scale in arcsec units based on the redshift and outflow size parameters (Section II.4).
@@
Figure 1: Example biconical outflow model from Bae and Woo (2016). The dust plane is shown as a red line between the two cones. The dark gray region represents the volume occupied by the outflowing gas (i.e., between inner and outer opening angles), while the light gray region denotes the hollow interior. The LOS direction is indicated by the black arrow. The definitions and adopted values of the cone and disk inclination angles are also shown.
II.2 Rotating Component
@@
While outflows are the dominant nongravitational component in AGN kinematics, the gas motions in the NLR are more complex due to additional dynamical effects. In particular, gas motion due to the gravitational potential of the host galaxy is represented by a core component of narrow emission lines, which show similar line width and offset compared to stellar absorption lines ($\sigma_{[\mathrm{OIII}],narrow}\sim\sigma_{*}$ Nelson and Whittle, 1996; Greene and Ho, 2005). If there is no nongravitational effect, gas kinematics are governed by a host galaxy potential, i.e., a rotating disk (e.g., Luo et al., 2019). Given that a majority of AGNs in the VVD diagram (see Figure 4(a)) show relatively small velocity offset and dispersion, the disk component should be included to reproduce the observed distribution of the gas kinematics properly. Even for AGNs with strong outflows, a rotational pattern is often observed in IFU data (Kang and Woo, 2018; Kim et al., 2023). Thus, a disk component is required to reproduce the characteristics in the spatially resolved kinematics.
@@
To generate disk models, we use the GalPak3D package (Galaxy Parameters and Kinematics, Bouché et al., 2015), which was originally developed for the morphokinematics of 3D observational data. Since our goal is to investigate the effect of a rotating disk on the observed gas kinematics, we employ only the basic models generated by the package without adjusting details. Using the default models of GalPak3D, we assume radial and perpendicular flux profiles as exponential ($n=1$) and Gaussian, respectively, and set the rotation curve to a hyperbolic tan (tanh) profile.
@@
We regulate two flux parameters ($f_{disk}/f_{cone}$ and $r_{1/2}$) and three kinematic parameters ($V_{max,rot}$, $r_{t}$, and $\sigma_{0}$) of the disk model. The inclination and PA of the disk are assumed to be the same as those of the dust plane in the outflow model. For the flux, the total disk flux within the SDSS fiber size (3″) is first determined by the disk-to-cone flux ratio parameter ($f_{disk}/f_{cone}$) and the integrated flux of the outflow. We then determine the disk flux distribution by assuming an exponential disk with a given half-light radius parameter ($r_{1/2}$).
@@
where $r$ is the radius on the galactic plane. At $r>r_{t}$, the rotational velocity begins to follow a flat radial profile. In addition, we adjust the intrinsic velocity dispersion ($\sigma_{0}$), which corresponds to the dynamical hotness of the disk.
@@
II.3 Seeing Effect
The beam smearing effect by atmospheric seeing has been pointed out as a significant obstacle in spatially resolved studies of AGNs (e.g., Husemann et al., 2016; Villar-Martín et al., 2016; Singha et al., 2022). Given that the seeing effect is more pronounced at higher redshifts, we need to include it in our model to account for outflows in a majority of AGN host galaxies.